OpenStudy (anonymous):

Your friend bought 3 bags of wild bird seed and an $18 bird feeder. Each bag of birdseed costs the same amount. If your friend spent $45, find the cost of one bag of birdseed.

OpenStudy (anonymous):

3x + 18 = 45 3x = 45 - 18 3x = 27 x = 27/3 x = 9 each bag of wild bird seed costs $9